what number can be subtracted from the numerator and denominator of the fraction 14/17 so that the value of the new fraction will 9/11?
Answer by checkley77(12844)

14/17-x=9/11
-x=9/11-14/17
-x=(9*17-14*11)/11*17
-x=(102-154)/187
-x=-52/187
-x=-52/187
x=52/187 ans.